OpenStudy (anonymous):

I need help finding an equation of the line that is parallel to the line 2x-3y=-4.

OpenStudy (anonymous):

Parallel lines have the same slope.

OpenStudy (anonymous):

2x-3y=4 -3y=4-2x y=(-3/4)+(2/3)x y=2/3x-.75 So, any line with a slope of 2/3

OpenStudy (anonymous):

thank you